1. An aqueous solution contains 0.395 M hydrocyanic acid. Calculate the pH of the solution after the addition of 5.99×10-2 moles of sodium hydroxide to 255 mL of this solution. (Assume that the volume does not change upon adding sodium hydroxide)

pH =

2. An aqueous solution contains 0.302 M acetic acid. Calculate the pH of the solution after the addition of 2.56×10-2 moles of potassium hydroxide to 125 mL of this solution. (Assume that the volume does not change upon adding potassium hydroxide)

pH =
Answer #1

1. no of moles of HCN = molarity * volume in L

                                  = 0.395*0.255   = 0.100725moles

Pka of HCN   = 9.31

------------- HCN(aq) + NaOH(aq) -------------> NaCN(aq) + H2O(l)

I------------0.100725 -----0.0599 ----------------------0

C--------- -0.00599 ------ -0.0599 -------------------- 0.0599

E------- 0.040825 ------- 0 -----------------------------0.0599

       PH   = PKa + log[NaCN]/[HCN]

              = 9.31 + log0.0599/0.040825

               = 9.31 + 0.1665

               = 9.4765 >>>>answer

2.

no of moles of CH3COOH = molarity * volume in L

                                  = 0.302*0.125 = 0.03775moles

Pka of CH3COOH = 4.74

------------- CH3COOH(aq) + KOH(aq) -------------> CH3COOKaq) + H2O(l)

I------------0.03775 -----   0.0256 --------------------------------0

C--------- -0.0256 ------ -0.0256 --------------------------- 0.0256

E------- 0.01215------- 0 ------------------------------------0.0256

       PH   = PKa + log[CH3COONa]/[CH3COOH]

              = 4.74 + log0.0256/0.01215

             = 4.74 + 0.3236

            = 5.0636 >>>>answer